find the inverse to the matrix M =
1 -1 2
-1 3 2
2 1 -3

and use this inverse to solve the system of equations
x-y+2z=1
-x+3y+2z=1
2x+y-3z=5
1
Expert's answer
2012-10-11T11:28:14-0400
The inverse matrix is

11/26 1/26 4/13
N^(-1) = -1/26 7/26 2/13
& 7/26 3/26& -1/13


By the way, you can watch our video explaining how to find inverse matrices



Let's solve the given system:

& & & & (1)
(x, y, z) = (1) * N^(-1) = (45/26, 23/26, 1/13).
& & & & & & & & (5)